OpenStudy (anonymous):

A circle is inscribed in a square. Write and simplify an expression for the ratio of the area of the square to the area of the circle. For a circle inscribed in a square, the diameter of the circle is equal to the side length of the square. Walk Me Through Please !?

OpenStudy (stamp):

I do not have time to consider the whole problem, but here is a picture. I would use a unit for the radius, and A=pi(r*r) and A = l*w 

OpenStudy (anonymous):